OpenStudy (anonymous):

Solve for r: |5r - 18| = 0. Show your work.

OpenStudy (turingtest):

answer or process?

OpenStudy (anonymous):

Well, both if you can.

OpenStudy (turingtest):

sure, look at it this way: the absolute value sign is meaningless, because if |a|=0 then a=0 that is the only way so you just need to solve 5r-18=0

OpenStudy (anonymous):

So would that be 3.6 or {3.6, -3.6}?

OpenStudy (turingtest):

5r-18=0 5r=18 r=18/5=3.6 since we dropped the absolute value signs there is no negative answer, just the one: 3.6

OpenStudy (anonymous):

Ok thanks